# Benchmarking and Building Zero-Shot Hindi Retrieval Model with Hindi-BEIR and NLLB-E5

### v2

@@ Abstract
Given the large number of Hindi speakers worldwide, there is a pressing need for robust and efficient information retrieval systems for Hindi. Despite ongoing research, comprehensive benchmarks for evaluating retrieval models in Hindi are lacking. To address this gap, we introduce the Hindi-BEIR benchmark, comprising 15 datasets across seven distinct tasks. We evaluate state-of-the-art multilingual retrieval models on the Hindi-BEIR benchmark, identifying task and domain-specific challenges that impact Hindi retrieval performance. Building on the insights from these results, we introduce NLLB-E5, a multilingual retrieval model that leverages a zero-shot approach to support Hindi without the need for Hindi training data. We believe our contributions, which include the release of the Hindi-BEIR benchmark and the NLLB-E5 model, will prove to be a valuable resource for researchers and promote advancements in multilingual retrieval models. The datasets from Hindi-BEIR are publicly available at Hindi-BEIR
@@ 1 Introduction
Information retrieval (IR) models are indispensable in our digital age, enabling swift and accurate extraction of relevant data from vast amounts of information. These models enable quick and accurate retrieval of relevant data, thereby facilitating informed decision-making across various downstream applications and domains. In this modern era of LLMs, retrievers have become even more relevant and almost necessary in curbing hallucinations from Large Language Model (LLM) generations by means of much popular retrieval augmented generation (RAG) Lewis et al. (2021) methods. However, the focus of information retrieval (IR) research has predominantly centered on the English language, which can be attributed to the availability of comprehensive benchmarks, such as the BEIR Thakur et al. (2021). Although, more recently, there have been efforts to build multi-lingual retrievers and associated evaluation benchmarks  Zhang et al. (2023, 2021) on non-English languages, multi-lingual expansion of robust benchmarks is still a work in progress, with many languages like Hindi lacking a BEIR like a comprehensive benchmark.
Hindi is one of the official languages of India [^1] https://en.wikipedia.org/wiki/Languages_of_India, a language of immense importance and global reach, with over half a billion speakers worldwide and ranking as the third most widely spoken language [^2] https://www.thecollector.com/what-are-the-most-spoken-languages-in-the-world/. However, from NLP research point of view, Hindi can be categorized as a low-resource language, lacking comprehensive resource and benchmarks to advance scientific research and this includes lack of a good retrieval benchmark too. Thus far, the evolution of Hindi or, in general, Indic language retrieval datasets has been rather ad-hoc, restricted to a small subset of domains, languages and tasks. Although a recent effort attempts to incorporate multiple Indic languages Haq et al. (2023) for a retrieval benchmark, the absence of diversity in domains and tasks remains a major limitation, affecting robust evaluation of the current state of retriever models in practical applications, which often involve varying tasks and domains.
@@
The introduction of Hindi-BEIR benchmark lets us assess and compare the existing multi-lingual retriever models more accurately for their zero-shot performance across diverse domains and tasks in Hindi-BEIR. Based on this performance comparison, we notice that a key limitation of existing models is the need for a substantial amount of language-specific training data, which is often not available for low-resource languages such as Hindi. To address this problem, we propose NLLB-E5, a multilingual retrieval model, inspired by recent works like Langbridge Yoon et al. (2024) and NLLB-LLM2Vec Schmidt et al. (2024), which distills information from a monolingual retrieval model into a multilingual retrieval model in a zero shot setup via a multi-lingual encoder. The key strength of NLLB-E5 is it does not need multilingual training data for distillation and thus eliminates the need for supervised multilingual training data.
@@ 2 Related Works
Our NLLB-E5 model is inspired by the works by Yoon et al. (2024) who propose Langbridge, an approach which enables multilingual reasoning without multilingual supervised training and Schmidt et al. (2024) who opt for a distillation based approach and introduce NLLB-LLM2Vec capable of producing robust multilingual embeddings extending from NLLB encoder.
@@ 3 Hindi-BEIR Retrieval Benchmark
MIRACL Question Answering 3494 350 66.49 10.42
@@
To ensure that the Hindi-BEIR Benchmark is comprehensive, challenging, and accessible to the public for future research and evaluation, we adhered to the following objectives:1) Diverse Domains and Tasks: We include datasets from diverse domains like Wikipedia and news articles to niche domains like scientific publications, finance to test the robustness and generalization ability of the retrieval models. Additionally, a good retrieval model should handle documents and queries of varying lengths equally well. We have ensured that the datasets included in the Hindi-BEIR benchmark exhibit a wide range of query and document lengths among the datasets. 2) Difficulty Level: We ensure that the datasets are challenging and systems relying solely on lexical overlap have a hard time retrieving the correct document. 3) Public Availability: All datasets curated in Hindi-BEIR Benchmark have friendly, user-friendly licenses and will be made publicly available.
@@ 3.1 Translating English Datasets to Hindi:
This strategy enables us to leverage the wealth of existing high-quality datasets in English while making them accessible and useful for Hindi language information retrieval tasks. 9 out of the 15 datasets, which include Arguana, FiQA-2018, TREC-COVID, SCIDOCS, SciFact, Touché-2020, NQ, FEVER, and Climate-FEVER were created by this method.
@@ 3.3 Subset from existing Multilingual Datasets
We also include 5 publicly available datasets, namely MIRACL (dev split)Zhang et al. (2021), MLDR Chen et al. (2024), mMarco Bonifacio et al. (2022), IndicQARetrieval (developed by modifying the IndicQA Dataset Doddapaneni et al. (2023)) and WikiPediaRetrieval  [^7] https://huggingface.co/datasets/ellamind/wikipedia-2023-11-retrieval-multilingual-queries.
@@ 4 NLLB-E5: multilingual retriever
4.1 LangBridge
Yoon et al. (2024) proposed an approach to equip existing LLMs with multilingual capability without any multilingual supervised data. They vertically stack a multilingual encoder and an LLM, adding a projection/alignment layer between them. The input first passes through the multilingual encoder, the representation from the multilingual encoder passes through the alignment layer and then the output from the alignment layer is passed as input to the LLM. They use English-labeled data to fine-tune the overall model and update only the embedding layers and the alignment layer. The model showed better multilingual capabilities.
4.2 NLLB-LLM2vec
Schmidt et al. (2024) extended the LLM2vec model BehnamGhader et al. (2024) by vertically stacking a NLLB model Team et al. (2022) and LLM2vec model. Similar to LangBridge Yoon et al. (2024), they add a projection/alignment layer between the two and add LoRA parameters on the LLM. Different from LangBridge, they have a teacher model and a student model. They use LLM2vec as the teacher model and the NLLB-LLM2vec as the student model. The input is passed through both the teacher and the student models. They minimize the mean squared error between the final representations obtained from both the teacher and the student model. They train the model in two stages, the first stage being general alignment between the teacher and the student model, and, the second stage being task alignment. They show that the NLLB-LLM2vec model is able to obtain superior performance on various natural language understanding tasks.
4.3 NLLB-E5
Figure 1 shows the architecture of NLLB-E5 with the different components and data flow between them. We build NLLB-E5 on the back of two components (1) A multilingual encoder NLLB and (2) a monolingual retriever E5. We bank on the multilingual encoder to project semantically similar sentences across languages to a shared representation space. We hypothesize that embeddings coming from such a language-agnostic shared representation space can act as a strong prior that can be easily aligned for producing multilingual task-specific embeddings. Moreover, given the multilingual nature of the embedding coming directly from the multilingual encoder, the aforementioned task-specific alignment is language agnostic and, therefore, can be done with task-specific data in a high-resource language, too.
@@ 5.1 Training Process
Get Student Model Embeddings: The input sentence is first tokenized using the NLLB tokenizer, producing a sequence of tokens $\{t_{1},t_{2},\dots,t_{m}\}$. The frozen multilingual encoder processes this token sequence, generating token-level representations in the multilingual semantic space $\{H_{1},H_{2},\dots,H_{m}\}$. To align with the input structure expected by the E5 model, the token sequence is modified by prepending a token embedding corresponding to either [CLS] query: (for queries) or [CLS] passage: (for corpus), obtained from the E5 embedding layer ($H_{\emph{prior}}$), along with the representation of [SEP]($H_{\emph{post}}$) which is appended at the end. This is projected through a learnable layer $W$ as $\{H^{\prime}_{1},H^{\prime}_{2},\dots,H^{\prime}_{m}\}=W\{H_{\emph{prior}},H_{1},H_{2},\dots,H_{m},H_{\emph{post}}\}$ This projected sequence is then passed through the student E5 model, which outputs the token-level representations $\{X^{\prime}_{1},X^{\prime}_{2},\dots,X^{\prime}_{m}\}$ Finally, average pooling is applied over these token-level representations to obtain the final sentence-level embedding $\vec{X}_{S}$ from the student model.
@@ 6.2 Implementation Details of NLLB-E5
To make NLLB-E5 produce robust generalizable embeddings across different kind of retrieval tasks and domains, we rely on using diverse data to learn NLLB and E5 alignment during training. We followed a similar approach as taken by generic text-to-embedding models and used a subset of data from the huge pool of  1B diverse data curated for training sentence transformer models SIMOULIN (2021). We randomly selected 100,000 samples from each dataset to create the final training set. The datasets chosen are i) Arxiv ii) Natural Questions (NQ) iii) HotpotQA iv) Stack Exchange - Duplicate questions (titles) v) Stack Exchange - (Title, Body) pairs vi) Stack Exchange - (Title+Body, Answer) pairs vii) Stack Exchange - (Title, Answer) pairs viii) Stack Exchange - Duplicate questions (bodies) ix) Stack Exchange duplicate questions (titles+bodies) x) SQuAD 2.0 xi) S2ORC Citation - pairs (Abstracts) xi) S2ORC (Title, Abstract) xii) SearchQA xiii) PubMed (Title, Abstract) xiv) SPECTER xv) FEVER xvi) Wikipedia Sections xvii) Stack Exchange Math xviii) Stack Overflow Posts xix) Wikipedia
@@ 7 Results and Analysis
In this section, we compare the performance of the aforementioned baselines with our proposed NLLB-E5 model on Hindi-BEIR benchmark and highlight our key findings from the results.
@@
Effect of Encoder Param Size: Table 3 presents the NDCG@10 scores for the NLLB model’s encoder across different parameter sizes. As the table shows, performance generally improves as the parameter size increases, though minor anomalies exist across specific datasets. Notably, the performance gain plateaus beyond the 1.3B parameter model, with only a marginal improvement of 0.09% when comparing the NLLB-E5 model’s 1.3B (distilled) encoder to that of the 3.3B parameter version. This possibly indicates that NLLB-E5 is not lacking in model capacity yet and thus can possibly be improved further by showing more volumes of data from diverse task and domains,
@@ 9 Limitations
While the Hindi-BEIR Benchmark and NLLB-E5 model provides valuable advancements, we acknowledge several limitations. One limitation of the Hindi-BEIR Benchmark is that it may not fully capture the breadth of tasks where retrieval models are essential. The current scope, while extensive, lacks coverage in critical domains such as Law and Medicine, which are key areas we plan to include in future expansions. Additionally, although designed primarily for Hindi, the benchmark’s extension to other languages is an area that remains unexplored in this work.
The NLLB-E5 model, while showing promise, exhibits sub-optimal performance on certain English tasks. This limitation indicates a need to explore alternative multilingual encoders to improve language performance. Additionally, we realise that the current model would architecture struggle with long texts due to a context length limitation of 508 tokens, which poses challenges in tasks requiring extended context handling. Furthermore, our approach to hyper-parameter tuning warrants further investigation to optimize model convergence and performance. Addressing these limitations will be crucial to the future development of more effective retrieval models and evaluation frameworks.
@@ A.2.3 TREC-COVID
An example of query with its corresponding golden corpus from the TREC-COVID dataset has been provided in Figure 8

### v3

@@ Abstract
Given the large number of Hindi speakers worldwide, there is a pressing need for robust and efficient information retrieval systems for Hindi. Despite ongoing research, comprehensive benchmarks for evaluating retrieval models in Hindi are lacking. To address this gap, we introduce the Hindi-BEIR benchmark, comprising 15 datasets across seven distinct tasks. We evaluate state-of-the-art multilingual retrieval models on the Hindi-BEIR benchmark, identifying task and domain-specific challenges that impact Hindi retrieval performance. Building on the insights from these results, we introduce NLLB-E5, a novel multilingual retrieval model that leverages a zero-shot approach to support Hindi without the need for Hindi training data. We believe our contributions, which include the release of the Hindi-BEIR benchmark and the NLLB-E5 model, will prove to be a valuable resource for researchers and promote advancements in multilingual retrieval models. The datasets from Hindi-BEIR are publicly available at Hindi-BEIR.The training and evaluation code for the NLLB-E5 model can be found on GitHub: NLLB-E5 Repository.
@@ 1 Introduction
Information retrieval (IR) models are indispensable in our digital age, enabling swift and accurate extraction of relevant data from vast amounts of information. These models enable quick and accurate retrieval of relevant data, thereby facilitating informed decision-making across various downstream applications and domains. In this modern era of LLMs, retrievers have become even more relevant and almost necessary in curbing hallucinations from Large Language Model (LLM) generations by means of much popular retrieval augmented generation (RAG) Lewis et al. (2021) methods. However, the focus of information retrieval (IR) research has predominantly centered on the English language, which can be attributed to the availability of comprehensive evaluation benchmarks, such as the BEIR Thakur et al. (2021). Although, more recently, there have been efforts to build multi-lingual retrievers and associated evaluation benchmarks  Zhang et al. (2023, 2021) on non-English languages, multi-lingual expansion of robust benchmarks is still a work in progress, with many languages like Hindi lacking a BEIR like a comprehensive benchmark.
Hindi is one of the official languages of India [^1] https://en.wikipedia.org/wiki/Languages_of_India, a language of immense importance and global reach, with over half a billion speakers worldwide and ranking as the third most widely spoken language [^2] https://www.thecollector.com/what-are-the-most-spoken-languages-in-the-world/. However, from NLP research point of view, Hindi can be categorized as a low-resource language, lacking comprehensive resource and benchmarks to advance scientific research and this includes lack of a good retrieval benchmark too. Though there has been some efforts in recent time for the development of Indic language dataset for various NLP task like Question answering (Sabane et al. (2024), Doddapaneni et al. (2023a)) and Summarization (Datta et al. (2023), Ghosh et al. (2024a), Ghosh et al. (2024b)), the evolution of Hindi or, in general, Indic language retrieval datasets has been rather ad-hoc, restricted to a small subset of domains, languages and tasks. Although a recent effort attempts to incorporate multiple Indic languages Haq et al. (2023) for a retrieval benchmark, the absence of diversity in domains and tasks remains a major limitation, affecting the robust evaluation of the current state-of-the-art retriever models in practical applications, which often involve varying tasks and domains.
@@
The introduction of Hindi-BEIR benchmark lets us assess and compare the existing multi-lingual retriever models more accurately for their zero-shot performance in Hindi across diverse domains and tasks covered in the benchmark. Based on this performance comparison, we notice that a key limitation of existing models is the need for a substantial amount of language-specific training data, which is often not available for low-resource languages such as Hindi. To address this problem, we propose NLLB-E5, a multilingual retrieval model, inspired by recent works like Langbridge Yoon et al. (2024) and NLLB-LLM2Vec Schmidt et al. (2024), which distills information from a monolingual retrieval model into a multilingual retrieval model in a zero shot setup via a multi-lingual encoder. The key strength of NLLB-E5 lies in its ability to perform distillation without reliance on multilingual training data, thereby eliminating the necessity for supervised multilingual datasets.
@@ 2 Related Works
Yoon et al. (2024) proposed an approach to equip existing LLMs with multilingual capability without any multilingual supervised data. They vertically stack a multilingual encoder and an LLM, adding a projection/alignment layer between them. The input first passes through the multilingual encoder, the representation from the multilingual encoder passes through the alignment layer and then the output from the alignment layer is passed as input to the LLM. They use English-labeled data to fine-tune the overall model and update only the embedding layers and the alignment layer. The model showed better multilingual capabilities. Schmidt et al. (2024) extended the LLM2vec model BehnamGhader et al. (2024) by vertically stacking a NLLB model Team et al. (2022) and LLM2vec model. Similar to LangBridge Yoon et al. (2024), they add a projection/alignment layer between the two and add LoRA parameters on the LLM. Different from LangBridge, they have a teacher model and a student model. They use LLM2vec as the teacher model and the NLLB-LLM2vec as the student model. The input is passed through both the teacher and the student models. They minimize the mean squared error between the final representations obtained from both the teacher and the student model. They train the model in two stages, the first stage being general alignment between the teacher and the student model, and, the second stage being task alignment. They show that the NLLB-LLM2vec model is able to obtain superior performance on various natural language understanding tasks.
Our NLLB-E5 model is inspired by Yoon et al. (2024) who propose LangBridge, an approach which enables multilingual reasoning without multilingual supervised training and Schmidt et al. (2024) who opt for a distillation based approach and introduce NLLB-LLM2Vec capable of producing robust multilingual embeddings extending from NLLB encoder.
@@ 3 Hindi-BEIR Retrieval Benchmark
MIRACL Passage Retrieval 506264 350 66.49 10.42
@@
To ensure that the Hindi-BEIR Benchmark is comprehensive, challenging, and accessible to the public for future research and evaluation, we adhered to the following objectives:1) Diverse Domains and Tasks: We include datasets from diverse domains like Wikipedia and news articles to niche domains like scientific publications, finance to test the robustness and generalization ability of the retrieval models. Additionally, a good retrieval model should handle documents and queries of varying lengths equally well. We have ensured that the datasets included in the Hindi-BEIR benchmark exhibit a wide range of query and document lengths among the datasets. 2) Difficulty Level: We ensure that the datasets are challenging and systems relying solely on lexical overlap have a hard time retrieving the correct document. 3) Public Availability: All datasets curated in Hindi-BEIR Benchmark have user-friendly licenses and are publicly available at Hindi-BEIR
@@ 3.1 Translating English Datasets to Hindi:
All entries that fell below this threshold were removed. If an entry was part of the corpus and had an associated query, the query was also discarded. Likewise, any corresponding records in the relevancy mapping file were also deleted.
This strategy enables us to leverage the wealth of existing high-quality datasets in English while making them accessible and useful for Hindi language information retrieval tasks. Nine out of the fifteen datasets, which include Arguana, FiQA-2018, TREC-COVID, SCIDOCS, SciFact, Touché-2020, NQ, FEVER, and Climate-FEVER were created by this method.
@@ 3.3 Subset from existing Multilingual Datasets
We also include 4 publicly available datasets, namely MIRACL (dev split)Zhang et al. (2021), mMarco Bonifacio et al. (2022), IndicQARetrieval (developed by modifying the IndicQA Dataset Doddapaneni et al. (2023b)) and WikiPediaRetrieval  [^7] https://huggingface.co/datasets/ellamind/wikipedia-2023-11-retrieval-multilingual-queries.
For translation, we specifically selected a subset of BEIR that focuses on factual and general knowledge content, avoiding topics with strong cultural or language-specific nuances that might be lost in translation. However, to ensure the inclusion of real-world data that captures the linguistic intricacies of Hindi, we carefully incorporate datasets such as Sangraha-IR, CC News Retrieval, MIRACL, IndicQA Retrieval, and Wikipedia Retrieval. These datasets, sourced from manually curated or web-scraped content, effectively preserve the cultural and language-specific nuances essential for robust evaluation.
@@ 4 NLLB-E5: multilingual retriever
Figure 1 shows the architecture of NLLB-E5 with the different components and data flow between them. We build NLLB-E5 on the back of two components (1) A multilingual encoder NLLB Team et al. (2022) and (2) a monolingual retriever E5 Wang et al. (2024a). We bank on the multilingual encoder to project semantically similar sentences across languages to a shared representation space. We hypothesize that embeddings coming from such a language-agnostic shared representation space can act as a strong prior that can be easily aligned for producing multilingual task-specific embeddings. Moreover, given the multilingual nature of the embedding coming directly from the multilingual encoder, the aforementioned task-specific alignment is language agnostic and, therefore, can be done with task-specific data in a high-resource language, too.
@@ 5.1 Training Process
Get Student Model Embeddings: The input sentence is tokenized using the NLLB tokenizer, producing a sequence of tokens $\{t_{1},t_{2},\dots,t_{m}\}$. The frozen multilingual encoder processes this token sequence, generating token-level representations in the multilingual semantic space $\{H_{1},H_{2},\dots,H_{m}\}$. To align with the input structure expected by the E5 model, the token sequence is modified by prepending a token embedding corresponding to either [CLS] query: (for queries) or [CLS] passage: (for corpus), obtained from the E5 embedding layer (referred from here on as $H_{\emph{prefix}}$), along with the representation of [SEP](referred from here on as $H_{\emph{postfix}}$) which is appended at the end. This is projected through a learnable layer $W$ as $\{H^{\prime}_{1},H^{\prime}_{2},\dots,H^{\prime}_{m}\}=W\{H_{\emph{prefix}},H_{1},H_{2},\dots,H_{m},H_{\emph{postfix}}\}$ This projected sequence is then passed through the student E5 model, which outputs the token-level representations $\{X^{\prime}_{1},X^{\prime}_{2},\dots,X^{\prime}_{m}\}$ Finally, average pooling is applied over these token-level representations to obtain the final sentence-level embedding $\vec{X}_{S}$ from the student model.
@@ 6.2 Implementation Details of NLLB-E5
To make NLLB-E5 produce robust, generalizable embeddings across different kind of retrieval tasks and domains, we rely on using diverse data to learn NLLB and E5 alignment during training. We followed a similar approach as taken by generic text-embedding models and used a subset of data from the huge pool of  1B diverse data curated for training sentence transformer models SIMOULIN (2021). We randomly selected 100,000 samples from each dataset to create the final training set. The datasets chosen are i) Arxiv ii) Natural Questions (NQ) iii) HotpotQA iv) Stack Exchange - Duplicate questions (titles) v) Stack Exchange - (Title, Body) pairs vi) Stack Exchange - (Title+Body, Answer) pairs vii) Stack Exchange - (Title, Answer) pairs viii) Stack Exchange - Duplicate questions (bodies) ix) Stack Exchange duplicate questions (titles+bodies) x) SQuAD 2.0 xi) S2ORC Citation - pairs (Abstracts) xi) S2ORC (Title, Abstract) xii) SearchQA xiii) PubMed (Title, Abstract) xiv) SPECTER xv) FEVER xvi) Wikipedia Sections xvii) Stack Exchange Math xviii) Stack Overflow Posts xix) Wikipedia
@@ 7 Results and Analysis
In this section, we compare the performance of the aforementioned baselines with our proposed NLLB-E5 model on Hindi-BEIR benchmark and highlight our key findings from the results. [^11] We also evaluated NLLB Encoder trained using distillation approach as used for NLLB-E5. Details has been discussed in A.6
@@
Effect of Encoder Param Size: Table 3 presents the NDCG@10 scores for the NLLB model’s encoder across different parameter sizes. As the table shows, performance generally improves as the parameter size increases, though minor anomalies exist across specific datasets. Notably, the performance gain plateaus beyond the 1.3B parameter model, with only a marginal improvement of 0.09% when comparing the NLLB-E5 model’s 1.3B (distilled) encoder to that of the 3.3B parameter version. This possibly indicates that NLLB-E5 is not lacking in model capacity yet and thus can possibly be improved further by showing more volumes of data from diverse tasks and domains.
@@ 9 Limitations
The queries are AI generated etc. While the Hindi-BEIR Benchmark and NLLB-E5 model provides valuable advancements, we acknowledge several limitations. One limitation of the Hindi-BEIR Benchmark is that it may not fully capture the breadth of tasks where retrieval models are essential. The current scope, while extensive, lacks coverage in critical domains such as Law and Medicine, which are key areas we plan to include in future expansions. Additionally, although designed primarily for Hindi, the benchmark’s extension to other languages is an area that remains unexplored in this work.
The NLLB-E5 model, while showing promise, exhibits sub-optimal performance on certain English tasks. This limitation indicates a need to explore alternative multilingual encoders to improve language performance. Additionally, we realise that the current model would architecture struggle with long texts due to a context length limitation of 508 tokens, which poses challenges in tasks requiring extended context handling. Further, we would also like to extend and validate this idea for other low-resource languages in the future.
@@ A.2.3 TREC-COVID
An example of a query with its corresponding golden corpus from the TREC-COVID dataset has been provided in Figure 8
@@ A.5 Ablation of LoRA parameters
A.6 Performance of NLLB Encoder
We trained the NLLB-Encoder using the same knowledge distillation approach as NLLB-E5 and evaluated its performance on subsets of both English and Hindi datasets from BEIR and Hindi-BEIR. The results, presented in Table 6, reveal a significant performance gap between NLLB-Encoder and both NLLB-E5 and mE5-Large in the retrieval task. These findings indicate that NLLB-Encoder alone is insufficient to act as a retrieval model and requires additional components, thereby motivating the development of NLLB-E5.
Model Hindi ArguAna Hindi FiQA-2018 IndicQARetrieval
NLLB(1.3B)-Encoder 44.45 17.54 42.98
NLLB-E5(1.3B) 57.24 34.19 61.94
mE5-Large 54.77 27.33 67.11
Table 6: NDCG@10 scores of NLLB-Encoder along with mE5-Large and NLLB-E5 on ArguAna and FiQA-2018 from Hindi-BEIR